# A Novel Shape-Aware Topological Representation for GPR Data with DNN Integration

### v2

@@ 4.4. Training Setup
  • Loss function: Binary cross entropy for classification and objectness, IoU loss for box regression (box = 0.05, cls = 0.5, iou_t = 0.2)
@@ 4.5. Results and Comparison
A predicted bounding box is considered correct if its Intersection over Union (IoU) with a ground truth box exceeds a specified threshold $0.6$. Matching is based on IoU overlap, and unmatched or duplicate predictions are penalized. This matching protocol is applied consistently across both simulated and real domains.
Table 4. Performance comparison across six configurations using 5-fold cross-validation (mean ± std). Results are reported separately for validation and test sets.
Model mAP@0.5 (Val) mAP@0.5 (Test) mAP@0.5:0.9 (Val) mAP@0.5:0.9 (Test) Precision (Val) Precision (Test) Recall (Val) Recall (Test)
Base 0.985±0.010 0.494±0.090 0.559±0.049 0.190±0.016 0.984±0.020 0.917±0.156 0.963±0.024 0.41±0.11
TE-Base 0.986±0.009 0.357±0.089 0.677±0.031 0.115±0.028 0.9649±0.037 0.787±0.220 0.9774±0.0201 0.293±0.035
@@
To better understand the individual contributions of the Sim2Real strategy and the proposed shape-aware topological representation, we conducted an ablation study focusing on Model Base, Model S2R, Model TE-Base, and Model TE-S2R. The results are illustrated in Fig. 5 (a) and Fig. 5 (b), corresponding to the validation and test sets, respectively.
@@
Among all configurations, Model TE-S2R which integrates both Sim2Real and shape-aware topological representation—consistently outperforms the others across all evaluation metrics. This confirms the effectiveness of the proposed approach in improving both accuracy and robustness, particularly under domain shift conditions.
@@ Conclusion
This paper presented a novel framework that combines shape-aware topological representations with deep learning-based object detection for GPR data. By incorporating lifetime-weighted persistent homology into the input representation, the proposed method enhances sensitivity to buried object signatures that may be obscured in conventional B-scan images. To the best of our knowledge, this is the first study to incorporated homological information as shape descriptors, marking a significant departure from conventional visual-only representations.
To address domain shift, a Sim2Real strategy was employed—pre-training on simulated data followed by fine-tuning on field data. Experimental results across multiple experiments configurations demonstrated that the TDA-enhanced YOLOv5 model consistently outperformed baseline models in terms of detection accuracy, recall, and precision.
The integration of topological descriptors contributes to both improved generalization and robustness, as these features encode global structural patterns that remain stable across varying environmental factors. This aligns with a well-known strength of topological data analysis (TDA), its ability to capture intrinsic shape information that remains invariant under noise, deformation, and other local perturbations. By relying on the topology of feature space rather than raw pixel intensities, the proposed method maintains stable performance even under challenging real-world GPR conditions.
These results suggest that topological priors can serve as a powerful complement to conventional visual features in GPR-based sensing. Furthermore, the framework may have broader implications beyond geophysical applications—such as biomedical imaging, where structural information is critical but often obscured by noise or intensity variations.

### v3

@@ 3.4. Training Setup
For the Sim2Real transfer-learning procedure, the YOLOv5s model was first pre-trained on 300 simulated B-scan images using standard detection losses (objectness, classification, and bounding-box regression). The learned weights from this pre-training stage were then transferred to the fine-tuning stage using real field data (Sites A and B), while the independent Site C dataset was reserved exclusively for testing. This process enables the model to retain fundamental reflection patterns learned from simulation and adapt them to domain-specific variations in real environments.
@@
  • Loss function: Binary cross entropy for classification and objectness, IoU loss for box regression (box = 0.05, cls = 0.6, iou_t = 0.2)
@@
Table 6. YOLOv5-based performance comparison across six model configurations using 5-fold cross-validation (mean ± std). Results are reported separately for validation and test sets.
Model mAP@0.5 (Val) mAP@0.5 (Test) mAP@0.5:0.95 (Val) mAP@0.5:0.95 (Test) Precision (Val) Precision (Test) Recall (Val) Recall (Test)
Base 0.986±0.009 0.357±0.089 0.677±0.031 0.115±0.028 0.9649±0.037 0.787±0.220 0.9774±0.0201 0.293±0.035
TE-Base 0.985±0.010 0.494±0.090 0.559±0.049 0.190±0.016 0.984±0.020 0.917±0.156 0.963±0.024 0.410±0.110
Table 7. YOLOv11-based performance comparison across six model configurations using 5-fold cross-validation (mean ± std). Only test-set results are reported to verify consistency across architectures.
Model mAP@0.5 (Test) mAP@0.5:0.95 (Test) Precision (Test) Recall (Test)
Base 0.773±0.108 0.326±0.040 0.845±0.126 0.65±0.067
Sim 0.040±0.089 0.016±0.036 0.050±0.112 0.017±0.037
S2R 0.795±0.094 0.353±0.087 0.873±0.086 0.703±0.088
TE-Base 0.771±0.046 0.360±0.040 0.894±0.061 0.714±0.050
TE-Sim 0.150±0.088 0.024±0.033 0.080±0.108 0.120±0.025
TE-S2R 0.806±0.016 0.378±0.024 0.934±0.052 0.748±0.004
@@ 3.5. Results and Comparison
To complement these numerical results, Fig.  5 visualizes the overall performance trends across all six configurations. The line chart clearly illustrates that model performance improves progressively from Sim to TE-S2R, demonstrating the consistent advantage of incorporating topological enhancement (TDA) and Sim2Real transfer learning. In particular, the proposed TE-S2R model achieves the highest Precision, Recall, and mAP values, confirming its superior accuracy and robustness under domain-shift conditions.
To further validate the robustness contribution of the proposed topological features, additional experiments were conducted using the YOLOv11 detector. To complement the numerical results, Fig.  5 compares the test-set performance trends of YOLOv5 and YOLOv11 across six configurations. This line chart highlights the consistency of model behavior between the two architectures: both exhibit progressive improvement from Base to TE-S2R, demonstrating the persistent benefit of incorporating topological enhancement (TDA) and Sim2Real transfer learning. In particular, the proposed TE-S2R configuration consistently achieves the highest Precision, Recall, and mAP values in both detectors, confirming its superior accuracy and robustness under domain-shift conditions.
The decrease in mAP and Recall on the test set mainly arises from the domain gap between the validation (Sites A and B) and test (Site C) datasets, which clearly demonstrates the independence of the test set. Nevertheless, Precision remains relatively stable, indicating that the model preserves discriminative reliability despite increased environmental variability. Across both YOLOv5 and YOLOv11 architectures, the TDA-enhanced configurations (TE-Base and TE-S2R) consistently exhibited higher stability and accuracy under noisy and heterogeneous soil conditions, particularly on the independent test dataset, further confirming that persistent homology improves global feature consistency and noise resistance.
To better understand the individual contributions of the Sim2Real strategy and the proposed shape-aware topological representation, we conducted an ablation study focusing on Model Base, Model S2R, Model TE-Base, and Model TE-S2R. The results are illustrated in Fig.  6 (a) and Fig.  6 (b), corresponding to the validation and test sets, respectively.
@@
Although the Base model trained purely on real data shows competitive results on the training sites (A and B), its performance does not generalize well to unseen domains due to the limited dataset size and site-specific conditions. The proposed Sim2Real and TDA-enhanced configurations mitigate this issue by pretraining on diverse simulated scenarios and transferring learned representations to real environments. This strategy significantly improves robustness and generalization, particularly on the independent Site C dataset.
Figure 5. Line chart comparing the test-set performance of YOLOv5- and YOLOv11-based models across six configurations.
@@
Among all configurations, Model TE-S2R which integrates both Sim2Real and shape-aware topological representation—consistently outperforms the others across all evaluation metrics. This confirms the effectiveness of the proposed approach in improving both accuracy and robustness, particularly under domain shift conditions. Notably, the TE-S2R model achieves higher Recall, mAP@0.5, and mAP@0.5:0.95 scores even when fine-tuned with the same limited amount of annotated field data, and further demonstrates improved generalization performance on the completely independent test set (Site C).
Figure 7. Qualitative comparison of detection results across representative configurations on a real B-scan image from the test site (Site C). Each greed bounding box indicates a detected buried pipe.
To visually demonstrate the effect of topological enhancement, Fig. 7 presents representative detection results on a real B-scan image from the independent test site (Site C). Compared with the Base and S2R configurations, the TDA-enhanced models (TE-Base and TE-S2R) produce clearer hyperbolic reflections and fewer false detections, visually confirming the quantitative improvements discussed above.
@@ Conclusion
This paper proposed a novel framework, TE-S2R (TDA-Enhanced Sim2Real), that integrates shape-aware topological representations with deep learning-based object detection for GPR data. By incorporating lifetime-weighted persistent homology into the input representation, the proposed method enhances sensitivity to buried object signatures that are often difficult to detect in conventional B-scan images. To the best of our knowledge, this is the first study to directly embed homological information as shape descriptors within a deep learning pipeline for GPR analysis, representing a meaningful departure from purely visual approaches.
A key contribution of this work lies in the synergistic integration of Topological Data Analysis (TDA) and Sim2Real (S2R) strategies. While S2R methods often suffer from domain gaps caused by structural inconsistencies and noise between synthetic and real-world data, TDA mitigates this issue by extracting global structural features that remain stable across domains. These topological features are inherently robust to local perturbations and help align the shape characteristics between simulated and field data. In TE-S2R, TDA not only strengthens the model’s resistance to noise but also enables a shared topological understanding across domains—transforming the S2R process from a simple domain adaptation step into a structure-aware transfer mechanism.
Experimental results demonstrate that TE-S2R consistently outperforms baseline models in terms of mean Average Precision (mAP), recall, and precision. The combination of topological priors and S2R enhances both generalization and robustness, enabling reliable detection even under real-world GPR conditions. These findings indicate that the integration of TDA and S2R is not merely additive but mutually reinforcing, yielding a performance gain that exceeds what either method can achieve alone.
Overall, the TE-S2R framework offers a scalable and noise-resistant solution for subsurface sensing tasks and can be readily extended beyond GPR to other imaging modalities where structural information is crucial but often obscured by noise or signal distortion.